A man is four times as old as his son.In five years time he wil be three times as old.what is their present ages
Answer by amarjeeth123(569)

let the son's age be x.
Then the man's age is 4x.
In five years time their ages will be (x+5) and (4x+5)
(4x+5)=3(x+5)
4x+5=3x+15
x=10
The man's age is 40.
The son's age is 10 and the man's age is 40 years.